At a pet shop there were 10 more dogs than cats. The shop sold 2 dogs and bought 2 cats. Now there are twice as many dogs than cats. How many cats does the shop have now?
Answer by macston(5194)

.
D=original number of dogs; C= original number of cats
.
D=C+10 Ten more dogs than cats
.
D-2=2(C+2) Substitute for C
(C+10)-2=2(C+2)
C+8=2C+4
4=C The shop originally had 4 cats.
The shop bought two cats so:
4+2=6
ANSWER: The shop now has 6 cats.
.
Original number of dogs:
D=C+10=4+10=14 The shop originally had 14 dogs.
The shop sold two dogs:
14-2=12 The shop now has 12 dogs.
12 dogs is twice as many as 6 cats.